OpenStudy (anonymous):

What integer needs to be added to the expression to create a perfect square? x^2 - 12x A. -36 B. 6 C. 24 D. 36

OpenStudy (lukecrayonz):

Take the second (12x) and divide by 2, then add to the other side. B. 6

OpenStudy (lukecrayonz):

WAIT no

OpenStudy (lukecrayonz):

take 12, divide by 2, then square it.

OpenStudy (lukecrayonz):

-12/2=-6^2=36

OpenStudy (anonymous):

Thank you so much.